OpenStudy (anonymous):

A bacteria culture grows with constant relative growth rate. After 2 hours there are 400 bacteria and after 8 hours the count is 50,000. Find the rate of growth after 7 hours. P'(7) = ?

OpenStudy (anonymous):

i was thinking to take the derivative of C*exp(r*7)

OpenStudy (anonymous):

the initial population of P(0) = 80

OpenStudy (anonymous):

C = 80

OpenStudy (anonymous):

i tried taking the natural log of 80*exp(r*7) and the taking log differentials

OpenStudy (paxpolaris):

\[\Large P(t)=C \cdot e^{r \cdot t}\] \[\large 400=Ce^{2r} \]\[\large 50000=Ce^{8r} \] \[\large\implies {500\cancel{00}\over 4\cancel{00}}={\cancel Ce^{8r}\over \cancel Ce^{2r}}\]\[\implies 125=e^{6r}\\\implies6r=\ln(125)=\ln \left( 5^3 \right)=3\ln(5)\\\implies \color{blue}{r={\ln \left( 5 \right)\over2}} \approx 0.804718956\] plugging this r in the first equation ... we can confirm \(C=80\)